(Consider using an Excel Spreadsheet to solve this problem.) Blueberry Farms Inc. is considering a four-year project to grow new and higher-quality blueberries. An initial investment of $20 million depreciable straight-line to zero over the projects life will buy the equipment necessary to get the project off the ground. The net working capital will also require an initial investment of $2.2 million to support the planting inventory; this cost is fully recoverable whenever the project ends. In the companys opinion, the project can generate pre-tax cash revenues of $19 million against the pre-tax cash operating costs of $6.2 million on an annual basis. The market value of the used equipment over the life of the project is as follows. At the End of Year 2 4 Market Value (S million $13.5 $11.7 $10.5 50.0 Consider the following: An income tax rate of 35% An appropriate discount rate of 15% All types of income are taxed at the same tax rate .If applicable, consider terminal loss or recapture based on salvage value a. If the company operates the project for four vears what is the NPV of the project? [6 points] b. If the company operates the project for: [9 points] a. b. c. the first vear only, what is the NPV of the project? the first two years only, what is the NPV of the project? the first three vears only, what is the NPV of the project? c. What economic life of the project maximizes its value to the company? If it is not four years then relative to operating the project for four years, what is the value of the option to abandon it after the value-maximizing economic life? [5 points]
Answer #1

Initial Investment = 20 + 2.2 = 22.2million
Pre tax profit = 19 - 6.2 - 5(depreciation) = 7.8 million
Tax Expense = 7.8*35% = 2.73 million
Therefore, Net cash flows per year = 19 - 6.2 - 2.73 = 10.07 million.

a) If co. operates for 4 years
Present value of cash inflows = 10.07*PVAF(4years,15%) + Terminal value * PVIF of 4th year
= (10.07 * 2.855) + 0 = 28.75 million
Initial Outlay = 22.2 million
Therefore, NPV = 28.75 - 22.2 = 6.55 million

b) a. If co. operates for 1year
  Present value of cash inflows = 10.07*PVAF(1years,15%) + Terminal value * PVIF of 1st year
= (10.07*0.8696) + 13.5*0.8696 = 20.4957
Therefore, NPV = 20.4957 - 22.2 = -1.7043 million

b. If co. operates for 2years
  Present value of cash inflows = 10.07*PVAF(2years,15%) + Terminal value * PVIF of 2nd year
= (10.07*1.6257) + 11.7*0.7561 = 25.2173
Therefore, NPV = 25.2173 - 22.2 = 3.0173 million

c. If co. operates for 3years
  Present value of cash inflows = 10.07*PVAF(3years,15%) + Terminal value * PVIF of 3rd year
= (10.07*2.2832) + 10.5*0.6575 = 29.8958
Therefore, NPV = 29.8958 - 22.2 = 7.6598 million

c) If the company operates the project for 3years, it maximises its value as the NPV is highest if it operates for 3 years. The value of the option is = 29.8958 million.